# A Look Into News Avoidance Through AWRS: An Avoidance-Aware Recommender System

### v1

@@ 5.4.2. Avoidance-aware Self-Attention Layer
where $\textbf{q}^{T}_{c}=\textbf{Q}_{c}\mathscr{N}$. This enhances the attention score with candidate-specific adjustments. The augmented representation $\textbf{l}^{k}_{i}$ for each click is formulated through attention weights:
(11) $\textbf{l}^{k}_{i}=\textbf{W}^{k}_{o}\sum_{j=1}^{N}\gamma^{k}_{j}\bm{h}_{j},\quad\gamma^{k}_{j}=\frac{\exp(r^{k}_{i,j})}{\sum_{p=1}^{N}\exp(r^{k}_{i,p})}$
@@
$\textbf{W}^{k}_{o}$ is the projection matrix for the $k$-th head. The comprehensive contextual representation $\textbf{l}_{i}$ for each click is derived by merging outputs from all $K$ attention heads. Note that by incorporating user engagement embeddings, our relatedness information now considers the influence of both avoidance and exposure, providing more context to user click patterns.
5.4.3. Avoidance-aware CNN Layer
The historical news vectors undergo processing through a convolutional neural network (CNN) equipped with self-attention mechanisms. According to (Qi et al., 2022), this involves applying multiple filters to capture patterns among local contexts of adjacent clicks and candidate news, formulated as
(12) $\textbf{s}_{i}=\textbf{W}_{awc}[\bm{h}_{i-h};\dots;\bm{h}_{i};\dots;\bm{h}_{i+h};\mathscr{N}]$
@@ 5.4.3. Avoidance-aware CNN Layer
5.4.4. Avoidance-aware Final Attention Layer
We use a candidate-aware attention network to model the importance of clicked news based on their relevance to avoidance-aware candidate news $\mathscr{N}$. This builds the avoidance-aware user embedding representation $\textbf{u}_{aw}$, given by $\textbf{u}_{aw}=\sum^{N}_{i=1}\alpha_{i}\textbf{m}_{i}$ where $\alpha_{i}$ is the weight of the $i$-th click:
(13) $\alpha_{i}=\frac{\exp(\Phi_{2}(\textbf{m}_{i},\mathscr{N}))}{\sum^{N}_{j=1}\exp(\Phi_{2}(\textbf{m}_{j},\mathscr{N}))}$
@@ 5.4.4. Avoidance-aware Final Attention Layer
5.4.5. Final Relevance Scores
@@
The relevance scores combine the news candidate’s avoidance-awareness $\mathscr{N}$ with the score from the avoidance-aware user vector $\textbf{u}_{aw}$. Specifically, the preliminary interest scores are calculated as $Int_{s}^{\prime}=\mathscr{N}^{T}\cdot\textbf{u}_{aw}$. However, to obtain the final scores, we apply our Avoidance-aware Relevance Predictor, which considers the time-varying nature of avoidance, as shown in Figure 4. Thus,
(14) $Int_{s}=(1-\eta)\cdot\mathbf{r}_{aw}+\eta\cdot Int_{s}^{\prime}$
@@
where $\eta=\sigma(\Phi_{3}(\mathbf{u}_{aw}))$. Here, $\Phi_{3}$ is a dense layer, $\mathbf{r}_{aw}$ is the relevance avoidance-aware score as defined earlier by equation 5, and $\eta$ is computed from the user representation $\mathbf{u}_{aw}$ using a dense network with sigmoid activation. This approach allows our model to better capture patterns related to the information conveyed by theavoidance and exposure values of a news article.
5.5. Loss Function
@@
Inspired by (Wu et al., 2019d), we use negative sampling during model training. For each clicked news item (positive sample), we randomly sample $K$ non-clicked items (negative samples) from the same impression and shuffle their order. The click probability score of the positive item is $\hat{y}^{+}$, and the scores for the $K$ negative items are $[\hat{y}^{-}_{1},\hat{y}^{-}_{2},\ldots,\hat{y}^{-}_{K}]$. Then, these scores are normalized using the softmax function given by
(15) $p_{i}=\frac{\exp(\hat{y}^{+}_{i})}{\exp(\hat{y}^{+}_{i})+\sum_{j=1}^{K}\exp(\hat{y}^{-}_{i,j})}$
We frame the click prediction task as a pseudo $(K+1)$-way classification problem. The loss function is the negative log-likelihood of all positive samples $S$: $\mathcal{L}=-\sum_{i\in S}\log(p_{i})$. This method trains the model to distinguish between clicked and non-clicked news items.
6. Experiment
6.1. Experimental Settings
6.1.1. Datasets
@@
We evaluated AWRS on three datasets: MIND-small (English) (Wu et al., 2020a), Adressa one-week (Norwegian) (Gulla et al., 2017), and Nikkei (Japanese). Nikkei Inc. [^2] https://www.nikkei.com/, based in Japan, is a leading media corporation renowned for its comprehensive coverage of business, economic, and financial news. Table 1 summarizes the train, validation, and test splits used in our study. For Nikkei, we analyzed 15,803 news articles, dividing user click behavior into training (Jan 16-20, 2023), validation (Jan 21, 2023), and testing (Jan 22, 2023) sets. The Adressa one-week dataset includes 22,136 news articles, with user click behaviors split into training (Jan 1-5, 2017), validation (Jan 6, 2017), and testing (Jan 7, 2017) sets. And the MIND-small dataset comprises 93,698 news articles, with user click behaviors segmented into training (Nov 9-13, 2019), validation (Nov 14, 2019), and testing (Nov 15, 2019) sets.
Table 1. Dataset Statistics

### v2

@@ Abstract
We adaptively select significant long-range relatedness metrics to model user interest in the candidate news $\bm{n}_{c}\in\mathscr{N}$ based on their contextual relevance $r^{k}_{i,j}=\hat{r}^{k}_{i,j}+\textbf{q}^{T}_{c}\textbf{W}^{r}_{k}\bm{h}_{j}$ where $\textbf{q}^{T}_{c}=\textbf{Q}_{c}\bm{n}_{c}$. This enhances the attention score with candidate-specific adjustments. The augmented representation $\textbf{l}^{k}_{i}$ for each click is formulated through attention weights:
(5.10) $\textbf{l}^{k}_{i}=\textbf{W}^{k}_{o}\sum_{j=1}^{N}\gamma^{k}_{j}\bm{h}_{j},\quad\gamma^{k}_{j}=\frac{\exp(r^{k}_{i,j})}{\sum_{p=1}^{N}\exp(r^{k}_{i,p})}$
@@
$\textbf{W}^{k}_{o}$ is the projection matrix for the $k$-th head. The comprehensive contextual representation $\textbf{l}_{i}$ for each click is derived by merging outputs from all $K$ attention heads.
Note that by incorporating user engagement embeddings, our relatedness information now considers the influence of both avoidance and exposure, providing more context to user click patterns.
5.4.3 Avoidance-aware CNN Layer.
@@
Table 1: Performance comparison for the Nikkei one-week, MIND-small, and Adressa one-week datasets. The best results are bolded, and the second best are underlined. All the results reported here used $D=5$.
Nikkei one-week MIND-small Adressa one-week
NRMS 51.87$\pm$2.3 30.55$\pm$1.9 20.42$\pm$2.7 29.99$\pm$2.5 49.64$\pm$0.7 24.37$\pm$4.4 22.20$\pm$4.7 28.80$\pm$4.8 55.65$\pm$5.1 21.88$\pm$1.6 16.49$\pm$3.1 26.11$\pm$5.2
NAML 50.11$\pm$0.0 33.23$\pm$1.0 23.24$\pm$0.8 33.22$\pm$0.9 48.54$\pm$2.5 26.88$\pm$4.1 24.84$\pm$4.4 31.17$\pm$4.4 50.00$\pm$0.0 23.52$\pm$0.7 21.81$\pm$1.7 29.43$\pm$0.0
LSTUR 50.36$\pm$1.2 29.03$\pm$3.6 19.61$\pm$4.1 29.29$\pm$4.1 50.01$\pm$0.0 29.31$\pm$1.2 27.50$\pm$1.4 33.95$\pm$1.3 54.87$\pm$3.2 32.54$\pm$2.8 30.31$\pm$3.2 36.92$\pm$3.0
TANR 50.02$\pm$0.0 29.15$\pm$0.8 20.39$\pm$0.2 29.81$\pm$0.3 52.24$\pm$1.5 29.41$\pm$1.9 27.73$\pm$1.5 34.25$\pm$1.4 52.39$\pm$4.1 24.25$\pm$4.4 22.65$\pm$5.0 29.67$\pm$4.2
SentiRec 50.00$\pm$0.0 29.46$\pm$0.9 20.51$\pm$0.1 29.96$\pm$0.0 52.89$\pm$2.9 27.73$\pm$1.3 26.42$\pm$1.3 32.81$\pm$1.3 52.44$\pm$1.4 20.24$\pm$1.2 16.80$\pm$1.7 26.85$\pm$1.5
MINER 55.71$\pm$1.1 31.73$\pm$1.1 23.04$\pm$0.3 34.65$\pm$1.3 61.41$\pm$0.7 27.26$\pm$0.5 25.91$\pm$0.7 32.74$\pm$0.5 54.43$\pm$5.2 26.40$\pm$7.2 22.55$\pm$11.2 31.41$\pm$11.0
MINS 50.68$\pm$1.3 27.78$\pm$1.0 17.21$\pm$0.8 26.58$\pm$0.3 47.60$\pm$4.3 24.72$\pm$4.3 23.00$\pm$4.8 29.28$\pm$4.8 69.88$\pm$0.5 35.34$\pm$0.4 36.35$\pm$3.5 44.35$\pm$4.6
CenNewsRec 52.68$\pm$0.2 31.20$\pm$2.1 20.97$\pm$3.6 29.22$\pm$3.5 49.46$\pm$0.9 20.88$\pm$0.8 18.59$\pm$0.8 24.92$\pm$0.9 53.99$\pm$5.1 28.86$\pm$2.9 31.46$\pm$3.9 36.32$\pm$1.5
MANNeR-CR 50.00$\pm$0.0 33.74$\pm$0.4 24.53$\pm$0.3 34.93$\pm$0.4 62.23$\pm$2.4 30.72$\pm$1.8 28.81$\pm$1.8 35.43$\pm$1.7 53.63$\pm$6.3 25.66$\pm$3.5 22.40$\pm$5.0 30.62$\pm$5.1
PP-REC 49.70$\pm$1.5 27.31$\pm$1.2 18.70$\pm$0.1 28.09$\pm$0.2 56.86$\pm$1.6 29.12$\pm$1.0 26.95$\pm$1.1 33.01$\pm$1.0 72.83$\pm$5.1 52.53$\pm$9.1 51.25$\pm$9.9 55.94$\pm$8.4
CAUM 55.27$\pm$2.2 32.89$\pm$2.2 23.51$\pm$2.8 32.76$\pm$3.5 55.76$\pm$0.7 31.68$\pm$0.6 30.04$\pm$0.6 36.40$\pm$0.6 60.87$\pm$4.4 25.66$\pm$1.8 25.20$\pm$2.2 31.74$\pm$6.0
AWRS 65.75$\pm$1.3 48.95$\pm$1.4 39.07$\pm$1.0 46.81$\pm$0.9 63.70$\pm$1.1 33.39$\pm$0.9 31.43$\pm$0.9 37.84$\pm$0.8 81.20$\pm$7.8 55.72$\pm$12.9 59.26$\pm$13.7 63.12$\pm$11.8
5.4.4 Avoidance-aware Final Attention Layer.
We use a candidate-aware attention network to model the importance of clicked news based on their relevance to avoidance-aware candidate news $\bm{n}_{c}\in\mathscr{N}$. This builds the avoidance-aware user embedding representation $\textbf{u}_{aw}$, given by $\textbf{u}_{aw}=\sum^{N}_{i=1}\alpha_{i}\textbf{m}_{i}$ where $\alpha_{i}$ is the weight of the $i$-th click:
(5.11) $\alpha_{i}=\frac{\exp(\Phi_{2}(\textbf{m}_{i},\bm{n}_{c}))}{\sum^{N}_{j=1}\exp(\Phi_{2}(\textbf{m}_{j},\bm{n}_{c}))}$
@@
5.4.5 Final Relevance Scores.
@@
The relevance scores combine the news candidate’s avoidance-awareness $\bm{n}_{c}\in\mathscr{N}$ with the score from the avoidance-aware user vector $\textbf{u}_{aw}$. Specifically, the preliminary interest scores are calculated as $Int_{s}^{\prime}=\bm{n}_{c}^{T}\cdot\textbf{u}_{aw}$. However, to obtain the final scores, we apply our Avoidance-aware Relevance Predictor, which considers the time-varying nature of avoidance, as shown in Figure 4. Thus,
(5.12) $Int_{s}=(1-\eta)\cdot\mathbf{r}_{aw}+\eta\cdot Int_{s}^{\prime}$
@@
where $\eta=\sigma(\Phi_{3}(\mathbf{u}_{aw}))$. Here, $\Phi_{3}$ is a dense layer, $\mathbf{r}_{aw}$ is the relevance avoidance-aware score as defined earlier by equation 5.5, and $\eta$ is computed from the user representation $\mathbf{u}_{aw}$ using a dense network with sigmoid activation. This approach allows our model to better capture patterns related to the information conveyed by the avoidance and exposure values of a news article.
5.5 Loss Function.
@@
Following [14], we employ negative sampling during training. For each clicked news (positive sample), $K$ non-clicked items (negative samples) are randomly drawn from the same impression and shuffled. The click probability of the positive item is $\hat{y}^{+}$, and the scores for the $K$ negatives are $[\hat{y}^{-}_{1},\hat{y}^{-}_{2},\ldots,\hat{y}^{-}_{K}]$. These scores are normalized using:
(5.13) $p_{i}=\frac{\exp(\hat{y}^{+}_{i})}{\exp(\hat{y}^{+}_{i})+\sum_{j=1}^{K}\exp(\hat{y}^{-}_{i,j})}$
This forms a pseudo $(K+1)$-way classification problem, where the loss function is the negative log-likelihood: $\mathcal{L}=-\sum_{i\in S}\log(p_{i})$.
6 Experiment.
6.1 Datasets.
@@
We evaluated AWRS on three datasets: MIND-small (English) [31], Adressa one-week (Norwegian) [30], and a proprietary Japanese dataset from Nikkei News [^2] https://www.nikkei.com/, based in Japan, is a leading media corporation renowned for its comprehensive coverage of business, economic, and financial news. For Nikkei one-week, we analyzed 15,803 news articles, split into training (Jan 16-20, 2023), validation (Jan 21, 2023), and testing (Jan 22, 2023). The Adressa one-week dataset includes 22,136 articles, with training (Jan 1-5, 2017), validation (Jan 6, 2017), and testing (Jan 7, 2017). The MIND-small dataset contains 93,698 articles, divided into training (Nov 9-13, 2019), validation (Nov 14, 2019), and testing (Nov 15, 2019). Additionally, Appendix A has details on the dataset splits.
6.2 Baselines and Evaluation Metrics.
We conducted a comparative analysis of the AWRS model against several state-of-the-art baseline models to evaluate its performance. The models compared include: (1) NRMS [14], (2) NAML [13], (3) LSTUR [15], (4) TANR [53], (5) SentiRec [23], (6) MINER [57], (7) MINS [56], (8) CenNewsRec [55], (9) MANNeR-CR [^3] For the Nikkei one-week dataset with the MANNeR-CR model, we reduced the token limit due to memory constraints. [24], (10) PP-REC [27], and (11) CAUM [22]. In accordance with prior research [22], we evaluate model performance using AUC, MRR, nDCG@5 and nDCG@10.
Table 2: Performance comparison on the MIND-small and Adressa one-week datasets using $D=5$ and word embeddings. The best results are bolded, and the second best are underlined.
Dataset Model AUC MRR nDCG@5 nDCG@10
MIND-small NRMS 55.71$\pm$2.37 29.24$\pm$2.32 27.30$\pm$2.21 33.70$\pm$2.02
NAML 50.12$\pm$0.09 33.88$\pm$1.94 32.17$\pm$1.70 38.43$\pm$1.69
LSTUR 54.52$\pm$0.58 32.09$\pm$0.28 30.14$\pm$0.36 36.60$\pm$0.29
TANR 57.66$\pm$3.26 32.57$\pm$0.15 30.68$\pm$0.25 36.95$\pm$0.11
SentiRec 54.03$\pm$0.47 30.51$\pm$0.07 28.23$\pm$0.02 34.78$\pm$0.06
CenNewsRec 53.40$\pm$1.44 26.01$\pm$0.32 24.61$\pm$0.25 31.32$\pm$0.19
CAUM 60.91$\pm$0.99 33.77$\pm$0.47 31.69$\pm$0.47 38.19$\pm$0.44
GLORY 52.15$\pm$0.13 29.88$\pm$0.22 32.25$\pm$0.17 38.48$\pm$0.08
AWRS 66.21$\pm$0.56 34.15$\pm$0.27 32.51$\pm$0.17 38.96$\pm$0.23
Adressa one-week NRMS 59.55$\pm$2.85 22.76$\pm$2.63 20.50$\pm$4.30 30.59$\pm$2.81
NAML 50.18$\pm$0.32 23.54$\pm$3.15 21.36$\pm$4.99 31.38$\pm$5.89
LSTUR 64.58$\pm$1.04 27.56$\pm$0.95 28.02$\pm$1.50 35.83$\pm$0.18
TANR 55.33$\pm$1.27 21.19$\pm$1.00 17.30$\pm$1.78 27.36$\pm$1.08
SentiRec 47.85$\pm$0.40 19.43$\pm$0.14 15.83$\pm$0.47 21.73$\pm$0.00
CenNewsRec 58.07$\pm$5.57 21.36$\pm$5.57 19.06$\pm$8.13 27.56$\pm$7.23
CAUM 64.73$\pm$3.86 24.85$\pm$3.17 23.22$\pm$4.89 33.21$\pm$4.57
GLORY 57.52$\pm$5.59 20.24$\pm$0.80 18.52$\pm$1.11 27.61$\pm$2.66
AWRS 67.71$\pm$7.88 34.74$\pm$4.18 35.55$\pm$5.17 41.62$\pm$6.80
6.3 Environment Configuration.
During training, we utilized mixed precision with the Adam optimizer [37]. The learning rates were set to 1e-5 for MIND-small and Nikkei one-week, and 1e-6 for Adressa one-week, with $k=4$ negative samples. Models were trained for a maximum of 10 epochs for MIND-small with early stopping configured to halt training after 3 epochs without improvement. For Nikkei one-week and Adressa one-week the models were trained A single NVIDIA A100 GPU was used. Our model is implemented based on the newsreclib  [^4] https://github.com/andreeaiana/newsreclib and is available as the awrs_recsys  [^5] https://github.com/toyolabo/awrs_recsys. Avoidance and user engagement embeddings were computed hourly for MIND-small, every two hours for Nikkei one-week, and every five hours for Adressa one-week. The reported results were obtained by running each model three times and calculating the average and standard deviation.
7 Results.
We proceed with a comprehensive evaluation of AWRS by addressing the following research questions (RQ).
7.1 RQ1 (Accuracy).
Does the proposed AWRS method outperform existing baselines?
Table 1 compares the performance of our proposed AWRS model with baseline models. As shown, AWRS achieves notable improvements across most key metrics, highlighting its enhanced awareness of avoidance and exposure patterns. For instance, a user’s choice to read an article about a niche topic like chess—often widely avoided—suggests a strong personal interest. In contrast, an article on the Oscars, which attracts broad attention, may be clicked due to its popularity rather than genuine interest. These patterns evolve over time; during a period of controversy, chess articles may gain traction, whereas interest in the Oscars may decline when far from the event’s date. This underscores the strong connection between avoidance and time.
The baseline models employ sophisticated architectures to capture user behavior effectively. However, by integrating user engagement embeddings with our Avoidance-aware User Encoder and Avoidance-aware Relevance Predictor, as illustrated in Figures 4 and 5, AWRS provides a deeper understanding of user preferences. This enables a more avoidance-aware contextual assessment of each article, capturing both user interests and avoidance behavior more effectively.
7.1.1 RQ2 (Performance without PLMs).
How does our model perform in the absence of pretrained language models (PLMs)?
We investigated how excluding pretrained language models (PLMs) in favor of word embeddings GloVe embeddings [58] for Egnlish and BPEmb [59] for Norwegian would affect our model’s performance. For evaluation, we utilized the same datasplit as described in Section 6.1. In this comparison, we included several baseline models for a more comprehensive evaluation: (1) NRMS [14], (2) NAML [13], (3) LSTUR [15], (4) TANR [53], (5) SentiRec [23], (6) CenNewsRec [55], (7) CAUM [22], and (8) GLORY [21]. Table 2 demonstrates that, overall, AWRS exhibits superior performance across all key metrics. It is important to address the slight performance differences between GLORY [21] and other models. These differences may arise from the variation in time splits used. We trained all models for 20 epochs with a early stopping configured to stop training after 5 epochs without improvement.
7.1.2 RQ3 (Effect of Each Component).
What is the impact of each AWRS component on accuracy scores?
To evaluate the influence of each component in our architecture, specifically the Avoidance-aware User Encoder and the Avoidance-aware Relevance Predictor, we conducted experiments by removing each component separately for all datasets. As shown in Table 3, each component performs worse individually than when combined. This occurs because the Avoidance-aware User Encoder captures historical patterns of user behavior regarding avoidance, such as a preference for more or less avoided news articles. Meanwhile, the Avoidance-aware Relevance Predictor understands the temporal and global impact of avoidance, adjusting relevance based on the timing of recommendations. This dual approach improves the model’s accuracy in predicting user preferences.
Table 3: Effect of the Avoidance-aware User Encoder, denoted as “rel,” and the Avoidance-aware Relevance Predictor, denoted as “avoid,” components on AWRS.
Dataset AWRS Model AUC MRR nDCG@5 nDCG@10
MIND-small only rel 58.18 32.44 30.83 36.90
only avoid 58.38 30.77 28.84 35.11
avoid + rel 63.70 33.39 31.43 37.84
Adressa one-week only rel 63.54 36.21 39.48 45.26
only avoid 54.23 25.64 22.04 24.99
avoid + rel 81.20 55.72 59.26 63.12
Nikkei one-week only rel 63.62 46.38 37.01 45.34
only avoid 65.22 48.89 37.84 45.68
avoid + rel 65.75 48.95 39.06 46.81
7.1.3 RQ4 (Effect of Varying $D$).
How does accuracy change with different values of $D$?
To assess the impact of the parameter $D$ on the AWRS model, we conducted experiments with values $D=5,7,10,15,$ and $20$. As shown in Table 4, adjusting $D$ significantly influences model performance, albeit with nuanced effects. For the Nikkei one-week dataset, the best results are observed at $D=5$ and $D=7$, while the Adressa one-week dataset achieves optimal performance at $D=5$. In the case of the MIND-small dataset, the highest AUC score is attained at $D=20$, MRR peaks at $D=10$, and both nDCG@5 and nDCG@10 reach their highest values at $D=15$. These findings highlight the importance of conducting a grid search to identify the most effective $D$ values for each dataset.
Table 4: AWRS performance for different $D$ values.
Dataset Metrics D = 5 D = 7 D = 10 D = 15 D = 20
MIND-small AUC 63.70 62.13 64.30 64.79 64.94
MRR 33.39 32.14 34.99 34.81 34.09
nDCG@5 31.43 30.49 32.85 32.89 32.26
nDCG@10 37.84 36.84 39.14 39.23 38.64
Adressa one-week AUC 81.20 65.61 69.97 64.43 76.52
MRR 55.72 43.10 45.98 41.09 50.55
nDCG@5 59.26 41.72 47.99 38.69 53.00
nDCG@10 63.12 46.70 50.62 44.59 60.33
Nikkei one-week AUC 65.75 64.66 62.52 62..83 64.24
MRR 48.95 48.04 41.68 44.97 44.94
nDCG@5 39.06 37.32 31.78 34.05 34.35
nDCG@10 46.81 45.31 40.96 42.79 43.30
8 Conclusion.
8.1 Architecture Choices
As discussed in Section 5.4, CAUM [22] was selected for its ability to capture relationships between candidate and previously viewed news. By incorporating user engagement embeddings through the Avoidance-aware User Encoder and the Avoidance-aware Relevance Predictor, AWRS enhances the original model by aligning content similarity, user preferences, and avoidance patterns. We believe that the user engagement embedding $\bm{ue}$ can be implemented on top of other models, presenting a promising avenue for future research. Our experiments indicate that CAUM and AWRS exhibit comparable training, validation, and testing times, with the primary drawback of the latter being increased dataset loading times due to the additional avoidance-related information.
8.2 Summary.
This study underscores the importance of avoidance awareness in enhancing the predictive accuracy of recommender systems. By analyzing timely avoidance behaviors, we gain deeper insights into user preferences, allowing us to incorporate this context into our models. Our AWRS model, integrating the Avoidance-aware Relevance Predictor and Avoidance-aware User Encoder modules, demonstrates strong performance across three real-world datasets, improving various benchmark metrics. This approach not only boosts accuracy but also lays the groundwork for future research on user behavior through avoidance.
Acknowledgments
This work is partially supported by “Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” in Japan (Project ID: jh241004), JSPS KAKENHI Grant Number 23K28098, and the Monbukagakusho: MEXT (Ministry of Education, Culture, Sports, Science and Technology - Japan) scholarship.
A Dataset Statistics
Table 5 presents detailed statistics on the number of “Impressions” (Denoted as Impr.) and “Users” for each dataset used in this study: MIND-small [31], Adressa one-week [30], and the proprietary dataset Nikkei one-week. Additionally, Table 6 summarizes the training, validation, and testing periods for each dataset.
Table 5: Dataset Statistics